rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#5311 — FS#9234 — Optical loop Lyon

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
We have a fibre down this morning in the Lyon optical loop, securisation has taken over. No outage expected. A technician is en route to carry out optic measurements.
Date:  Monday, 02 September 2013, 20:04PM
Reason for closing:  Done
Comment by OVH - Monday, 02 September 2013, 14:26PM

We have carried out the measures on our side. We are no longer getting a signal on an optic strand between SPR and MPL. A ticket has been opened with our fibre provider.


Comment by OVH - Monday, 02 September 2013, 18:10PM

The cut in the fibre of the operator has been detected 2500m from the SPR site, a team has been deployed to takae care of the soldering.


Comment by OVH - Monday, 02 September 2013, 20:03PM

There was a broken fibre in an optic case. Everything is operational again, securisation is in place.